# MEGANet-W: A Wavelet-Driven Edge-Guided Attention Framework for Weak Boundary Polyp Detection

### v1

@@ Abstract
Colorectal polyp segmentation is critical for early detection of colorectal cancer, yet weak and low contrast boundaries significantly limit automated accuracy. Existing deep models either blur fine edge details or rely on handcrafted filters that perform poorly under variable imaging conditions. We propose MEGANet-W, a Wavelet Driven Edge Guided Attention Network that injects directional, parameter free Haar wavelet edge maps into each decoder stage to recalibrate semantic features. Our two main contributions are: (1) a two-level Haar wavelet head for multi orientation edge extraction; and (2) Wavelet Edge Guided Attention (W-EGA) modules that fuse wavelet cues with reverse and input branches. On five public polyp datasets, MEGANetW consistently outperforms existing methods, improving mIoU by up to 2.3% and mDice by 1.2%, while introducing no additional learnable parameters.
@@ 1 Introduction
U-Net along with variants, which are based on convolutional neural networks, have facilitated the rapid development of medical image segmentation tasks [29]. Subsequent studies use Transformer as encoder to strengthen the network’s global feature representation [32, 37]. Attention mechanism SE-Net proposed by [9] and CBAM introduced by [30] enhanced global feature selectivity and emphasizing salient regions. However, these models offen suffer from spatial down-sampling, leading to blurred boundaries, especially in low-contrast regions where edge features are critical.
To address edge preservation, recent works incorporate frequency-domain processing. [25] introduced FAENet, which apply discrete wavelet transform (DWT) to decompose features into high- and low-frequency bands, using both intra- and cross-band attention to sharpen edges and maintain internal texture consistency in polyp images. [39] proposed WRANet, which replaces pooling layers with DWT to better separate noise-prone high-frequency and stable low-frequency components. Furthermore, boundary-aware dynamic architecture such as WBANet, proposed by [29], integrate multi-scale feature fusion with wavelet-driven attention to reinforce the modelling of ambiguous polyp edges.
Beyond wavelet-based models, [4] introduced SFA, which enhances boundary modelling through an auxiliary decoder dedicated to edge prediction and a boundary-sensitive loss that enforces area-boundary consistency. [6] proposed PraNet, introducing reverse attention to improve segmentation in ambiguous boundary regions. Following this, MEGANet [2] preserve high-frequency edge cues by using Laplacian-derived map with Multi-Scale Edge-Guided Attention (EGA). However, it relies on classical edge detection and lacks deeper integration of frequency-domain embedding.
@@
Wavelet Edge‐Guided Attention modules: Building upon the Edge‐Attention module in MEGANet [2], we introduce a wavelet‐based extension in which edge cues derived from wavelet transforms are used at each decoder stage to dynamically recalibrate semantic feature maps confidence.
@@ 2 Related Work
We categorize relevant work into three domians: edge-aware segmentation, wavelet-based feature extraction, and attention mechanism. Our method builds on these by integrating wavelet-derived edge features into a decoder attention framework for improved polyp segmentation.
Edge-Aware Polyp Segmentation. Accurate edge preservation is crucial in polyp segmentation due to the subtle yet critical boundaries between polyps and surrounding mucosa [3]. Early approaches utilized hand-crafted edge detectors such as Canny, Sobel or Laplacian filters, but these suffered from sensitivity to noise and lacked semantic understanding [10]. Recent deep learning-based approaches integrating boundary awareness directly into segmentation networks [33]. [4] introduced SFA, which enhances boundary modelling through an auxiliary decoder dedicated to edge prediction and a boundary-sensitive loss that enforce area-boundary consistency. [6] proposed PraNet, introducing reverse attention to enhance segmentation in ambiguous boundary regions. Following this, [2] presented MEGANet, which combines encoder features, decoder predictions, and Laplacian-derived high-frequency maps via a Multi-Scale Edge-Guided Attention (EGA). [14] further refined boundaries by combining shallow features with reverse attention, while [15] and [36] incorporated dual attention mechanisms tailored for edge modeling. However, these approaches either rely on learned edge labels or hand-crafted filters, and do not leverage the frequency domain explicitly for boundary enhancement.
@@
Attention Mechanisms in Polyp Segmentation. Attention mechanisms have become foundational in modern segmentation architectures by enabling adaptive focus on informative spatial and channel-wise features. Initial efforts such as SE-Net proposed by [9] and CBAM introduced by [30] enhanced global feature selectivity, while more recent advances have targeted boundary precision in medical image segmentation. MSNet, proposed by [38], introduces a multi-scale subtraction module to suppress redundant information and extract complementary features, thereby improving attention-based segmentation across scales. [22] proposed Polyper, which introduces boundary-sensitive attention to distinguish polyp interiors from edges, while [35] presented LDNet, which employ lesion-aware cross-attention to refine ambiguous contours. [19] introduced PEFNet, which integrates positional encoding of polyp regions into skip connections, guiding attention to more salient spatial regions with structural priors. MISNet proposed by [13] employed a parallel attention structure with a balancing module to enhance boundary localization across scales.
Meanwhile, frequency-aware attention has gained traction for its ability to separate edge and context features. [25] introduced FAENet, which apply intra- and cross- frequency attention using Fourier filtering to preserve both local and global structure, and [7] designed WLAM, which combined wavelet-derived low- and high- frequency features within a linear attention framework. Although promising, these designs are often limited to encoder-side processing or classification tasks, and few apply frequency-domain signals directly within decoder attention. Most closely related, WBANet proposed by [29] utilized wavelet-based attention into a boundary-aware segmentation framework, yet lacks explicit edge-map injection into attention queries. Our work bridges this gap by feeding wavelet-derived edge features into decoder attention, enabling the model to adaptively focus on high-frequency boundary information during reconstruction, combining the strengths of spatial attention with the domain-adaptivity of wavelets.
@@ 3 Proposed Methodology
This section details the upgraded MEGANet-W, as despited in Fig.2. Relative to the original MEGANet baseline [2], we (i) keep the strong ResNet-34 encoder-decoder backbone, (ii) replace the Laplacian-based edge extractor with a two-level Haar wavelet head, and (iii) integrate the resulting multi-orientation edge cues into the decoder through an Wavelet Edge-Guided Attention (W-EGA) block.
@@ 3.1 Network Architecture
Our proposed network preserves the encoder–decoder macro‐structure of the baseline while threading edge information through every decoding step. Given an RGB image , a ResNet‐34 backbone pretrained on ImageNet produces five spatial‐scale feature maps at strides , respectively. Starting from , a symmetric decoder comprising five upsampling blocks restores the original resolution. At every scale , except the coarsest, we insert the W-EGA module (see Sec. 2), replacing the Laplacian‐based EGA used in the MEGANet baseline. Each decoder stage is projected to a side output via a Conv–BN–ReLU block, followed by a Conv prediction head. The side logits are bilinearly resized to the input resolution and jointly supervised. Table 1 summarizes the feature tensor dimensions across MEGANet-W.
Stage | Operator (stride) | Tensor shape () |
Conv7, 64 (2) | ||
MaxPool (2) + Res-34 Layer1 | ||
Res-34 Layer2 | ||
Res-34 Layer3 | ||
Res-34 Layer4 | ||
Conv | ||
Up (768 256) | ||
Up (384 128) | ||
Up (192 64) | ||
Up (128 64) |
@@ 3.2 Wavelet Enhanced Edge Head
This tensor serves as input to the subsequent W-EGA modules. The orthonormality of the Haar kernels ensures energy preservation while extracting directional edge bands. This head operates end-to-end with three fixed 2 × 2 filters, adds no learnable parameters, and introduces high-frequency priors absent from conventional CNN features. The resulting multi-scale edge representation is richer, enable it effectively encode structural boundaries across multiple resolutions.
@@ Acknowledgements
We appreciate Xin Ying, Zhe Enn, and Zhe Yang for providing the computational units used in this work. We also thank Leon Loo and Zheng Yeh for lending us their institutional accounts, which enabled us access to advanced academic materials.

### v2

@@ Abstract
Colorectal polyp segmentation is critical for early detection of colorectal cancer, yet weak and low contrast boundaries significantly limit automated accuracy. Existing deep models either blur fine edge details or rely on handcrafted filters that perform poorly under variable imaging conditions. We propose MEGANet-W, a Wavelet Driven Edge Guided Attention Network that injects directional, parameter free Haar wavelet edge maps into each decoder stage to recalibrate semantic features. Our two main contributions are: (1) a two-level Haar wavelet head for multi orientation edge extraction; and (2) Wavelet Edge Guided Attention (W-EGA) modules that fuse wavelet cues with boundary and input branches. On five public polyp datasets, MEGANet-W consistently outperforms existing methods, improving mIoU by up to 2.3% and mDice by 1.2%, while introducing no additional learnable parameters.
@@ 1 Introduction
U-Net along with variants, which are based on convolutional neural networks, have facilitated the rapid development of medical image segmentation tasks [29]. Subsequent studies use Transformer as encoder to strengthen the network’s global feature representation [32, 37]. Attention mechanism like SE-Net [9] and CBAM citeWoo2018 improve feature selectivity but offen blur boundaries due to spatial down-sampling, especially in low-contrast regions.
To enhance edge preservation, recent approaches leverage frequency-domain processing via discrete wavelet transform (DWT). Methods such as FAENet [25], WRANet [39], and WBANet [29] integrate wavelet-based decomposition and attention mechanisms to improve edge sharpness and texture consistency, particularly in polyp segmentation. Beyond wavelet-based models, SFA [4] uses boundary-aware supervision for edge refinement, while PraNet [6] applies reverse attention for ambiguous regions. MEGANet [2] introduces Laplacian-based edge cues but relies on classical edge detection and lacks deep frequency-domain integration.
@@
Wavelet Edge‐Guided Attention modules: A wavelet‐based extension uses edge cues derived from wavelet transforms at each decoder stage to recalibrate semantic feature maps confidence.
@@ 2 Related Work
We categorize relevant work into three domains: edge-aware segmentation, wavelet-based feature extraction, and attention mechanisms. Our method builds upon these by integrating wavelet-derived edge features within a decoder attention framework for improved polyp segmentation.
Edge-Aware Polyp Segmentation. Accurate edge preservation is crucial in polyp segmentation due to the subtle yet critical boundaries between polyps and surrounding mucosa [3]. Early approaches utilized hand-crafted edge detectors such as Canny, Sobel or Laplacian filters, but these suffered from sensitivity to noise and lacked semantic understanding [10]. Recent deep learning-based approaches integrating boundary awareness directly into segmentation networks [33]. SFA [4] enhances boundary modelling through an auxiliary decoder dedicated to edge prediction and a boundary-sensitive loss that enforce area-boundary consistency. PraNet [6], introduces reverse attention to enhance segmentation in ambiguous boundary regions. Following this, MEGANet [2] combines encoder features, decoder predictions, and Laplacian-derived high-frequency maps via a Multi-Scale Edge-Guided Attention (EGA). [14] further refined boundaries by combining shallow features with reverse attention, while [15] and [36] incorporated dual attention mechanisms tailored for edge modeling. However, these approaches either rely on learned edge labels or hand-crafted filters, and do not leverage the frequency domain explicitly for boundary enhancement.
@@
Attention Mechanisms in Polyp Segmentation. Attention mechanisms are foundational in modern segmentation architectures, enabling adaptive focus on informative spatial and channel-wise features. Early approaches such as SE-Net [9] and CBAM [30] enhanced global feature selectivity. More recent advances have targeted boundary precision in medical image segmentation. MSNet [38] introduces a multi-scale subtraction module to suppress redundant information and extract complementary features, improving attention-based segmentation across scales. Polyper [22] introduces boundary-sensitive attention to distinguish polyp interiors from edges, while LDNet [35] employs lesion-aware cross-attention to refine ambiguous contours. PEFNet [19] integrates positional encoding of polyp regions into skip connections, guiding attention toward structurally salient regions. MISNet [13] adopts a parallel attention structure with a balancing module to enhance boundary localization across scales.
Meanwhile, frequency-aware attention has gained traction for its ability to separate edge and context features. FAENet [25] apply discrete wavelet transform (DWT) to decompose features into high- and low-frequency bands, and WLAM [7] combines wavelet-derived low- and high- frequency features within a linear attention framework. Although promising, these designs are often limited to encoder-side processing or classification tasks, and few apply frequency-domain signals directly within decoder attention. Most closely related, WBANet [29] utilizes wavelet-based attention into a boundary-aware segmentation framework, yet lacks explicit edge-map injection into attention queries. Our work bridges this gap by feeding wavelet-derived edge features into decoder attention, enabling the model to adaptively focus on high-frequency boundary information during reconstruction, combining the strengths of spatial attention with the domain-adaptivity of wavelets.
@@ 3 Proposed Methodology
This section describes the upgraded MEGANet-W, as depicted in Fig.2. Compared to the original MEGANet-R34 baseline [2], the proposed model: (i) replaces the Laplacian-based edge extractor with a two-level Haar wavelet head, and (ii) integrates the resulting multi-orientation edge features into the decoder via a Wavelet Edge-Guided Attention (W-EGA) block.
@@ 3.1 Network Architecture
Our proposed network retains the encoder–decoder architecture of the [2] while integrating edge information at each decoding stage. Given an RGB image , a ResNet‐34 backbone extracts five spatial‐scale feature maps at strides , respectively. Starting from , a symmetric decoder with five upsampling blocks restores the input resolution. At each decoding stage , except the coarsest, we insert W-EGA module (see Fig. 2), replacing the Laplacian‐based EGA from [2]. Each decoder stage is projected to a side output via a standard ConvBlock (see Sec. 3.3), followed by a convolution. The side logits are bilinearly upsampled to the input resolution and supervised jointly.
@@ 3.2 Wavelet Enhanced Edge Head
This tensor is fed into the subsequent W-EGA modules. The orthonormality of the Haar kernels ensures energy preservation while extracting directional edge bands. This wavelet head operates end-to-end using three fixed 2 × 2 filters, introduces no learnable parameters, and embeds high-frequency priors typically absent from conventional CNN features. The resulting multi-scale edge representation is richer and effectively encodes structural boundaries across multiple resolutions.
@@ Acknowledgements
We appreciate Xin Ying, Zhe Enn, and Zhe Yang for providing the computational units used in this work. We also thank Leon Loo and Zheng Yeh for their support.